OpenStudy (anonymous):

can you graph |1-x|/1-x?

OpenStudy (anonymous):

Yes, depending on what the answer needs, simply plug in a few values of x and plot the values. For example at x=0 we have |1-0|/1-0 or 1/1, so you would plot the point (0,1)

OpenStudy (anonymous):

this is either -1 or 1 depending on whether x>1 or x< 1

OpenStudy (anonymous):



OpenStudy (anonymous):

\[|1-x| = \left\{\begin{array}{rcc} 1-x & \text{if} & x <1 \\ x-1& \text{if} & x >1 \end{array} \right. \] therefore \[\frac{|1-x|}{1-x} = \left\{\begin{array}{rcc} 1& \text{if} & x <1 \\ -1& \text{if} & x >1 \end{array} \right. \]